OpenStudy (anonymous):

After the softball game, Alise's coach bought each of the 15 players a double-scoop or triple-scoop ice cream cone at the local ice cream shop. The coach bought a total of 36 scoops of ice cream. Use a model to determine how many girls got a triple-scoop ice cream cone. 6 7 8 9

OpenStudy (anonymous):

The answer is 11 not any of those and I have proof.

OpenStudy (anonymous):

let x be the no of girls who took triple-scoop, so (15-x) girls took double-scoop therefore 3x + 2(15 - x) = total no of scoops => 3x + 2(15 - x) = 36 => x = 6 therefore 6 girls got a triple scoop

OpenStudy (anonymous):

It's 11 I can prove it.

OpenStudy (anonymous):

@Crazykid95 : i want to see your proof

OpenStudy (anonymous):

x + y = 15 x is number of girls who got double, y is number of girls who got triple. 2x + 3y = 36 Systems of equations. x = 15 - y 2(15-y) + 3y = 36 30 - 2y + 3y = 36 30 + y =36 y = 6 6 girls got triple, 9 girls got double.